# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 4.1 The figure-eight knot
$D^{({\rm conj})}_{-1}(q)=q\,,\quad D^{({\rm conj})}_{-s}(q)=0\quad s\geq 2\,.$ (4.143)
We have verified the integrality structure numerically up to $s=20$. It is then straightforward to compute the integers $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, using Eqs. (3.65), (3.66), which give
$\{c^{(\rm{geom})}_{s}\}_{s\in{\mathbb{Z}}_{>0}}$ =$=\{1,5,10,21,51,122,295,725,1792,4455,11133,27930,\dots\}\,,$ (4.144a)
$\{c^{(\rm{conj})}_{s}\}_{s\in{\mathbb{Z}}_{>0}}$ $=(-1)\{0,4,9,20,50,121,294,724,1791,4454,11132,27929,\dots\}\,,$ (4.144b)
while $c^{(\rm{conj})}_{-s}=1$ for all $s\in{\mathbb{Z}}_{>0}$. We point out that the integers in Eq. (4.144a) appear to match numerically the sequence [^11] This is the entry A132461132461132461132461 in the On-Line Encyclopedia of Integer Sequences.
$\sum_{k=0}^{\left\lfloor s/2\right\rfloor}\left({s-k\choose k}+{s-k-1\choose k-1}\right)^{2}\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (4.145)
@@
$c^{(\rm{geom})}_{s}+c^{(\rm{conj})}_{s}=c^{(\rm{conj})}_{-s}\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (4.146)
@@
Let us now go back to the quantum $A$-polynomial for the $\bm{4}_{1}$-knot written in Eq. (2.31). The corresponding second-order $q$-difference equation is
$\hat{A}_{\bm{4}_{1}}(\hat{P},\hat{X},q)\chi^{(\alpha),\rm{WKB}}(u;\hbar)=0\,,$ (4.147)
where $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ is the resummed perturbative wave function in Eq. (3.87) associated with the branch $\alpha$. In the rational case of $q={\rm e}^{2\pi{\rm i}{P\over Q}}$ with $P,Q\in{\mathbb{Z}}_{>0}$ coprime, Eq. (4.147) can be solved in closed form by directly applying the results of Section 3.4.1. For simplicity, let us consider the classical solution $\alpha=\rm{geom}$ and drop the label from our notation. After suitably taking into account the exponential prefactor containing the functions $s_{0,1}(x)$ in Eq. (4.138a), the $q$-difference equation above assumes the form in Eq. (3.90) with the identifications
$A(X)=-q^{-2}X^{-2}{X^{2}-1\over q^{-1}X^{2}-1}\tilde{a}(X;q)\,,\quad B(X)=q^{-2}X^{-4}{qX^{2}-1\over q^{-1}X^{2}-1}\,,$ (4.148)
where we have introduced the function
@@
Substituting the expressions in Eq. (4.148) into Eq. (3.99), we compute the matrix ${\cal M}_{Q}(X)$ and obtain explicit formulae for $\rm{tr}{\cal M}_{Q}(X)$, $\det{\cal M}_{Q}(X)$, and $\Delta_{Q}(X)$. Specifically, we find that
$\rm{tr}{\cal M}_{Q}(X)$ $=X^{-4Q}\left(1-X^{Q}-2X^{2Q}-X^{3Q}+X^{4Q}\right)\,,$ (4.150a)
@@
$\Delta_{Q}(X)$ $=X^{-8Q}\left(1-X^{2Q}\right)^{2}\left(1-2X^{Q}-X^{2Q}-2X^{3Q}+X^{4Q}\right)\,,$ (4.150c)
and plugging these into Eq. (3.4.1) with a choice of plus sign to satisfy the requirement in Eq. (3.109) for $\alpha=\text{geom}$, we obtain the functions $\Upsilon(X)$ and $\varphi_{Q}(X)$. Observe that all functions in Eq. (4.1), and thus also $\varphi_{Q}(X)$, depend on $X$ through $X^{Q}$, as expected, and we can identify
$\rm{tr}{\cal M}_{Q}(X)$ $=X^{-4Q}T(X^{Q})=X^{-2Q}\tilde{a}(X^{Q};1)\,,$ (4.151a)
$\Delta_{Q}(X)$ $=X^{-8Q}\left(1-X^{2Q}\right)^{2}\Delta(X^{Q})=X^{-4Q}\left(\tilde{a}(X^{Q};1)^{2}-4\right)\,,$ (4.151b)
where $\tilde{a}(X;q)$ is defined in Eq. (4.149), while $T(X)$ and $\Delta(X)$ are the functions introduced in Eqs. (2.28a), (2.28b), respectively, which appear in the formula for the solutions to the classical $A$-polynomial of the figure-eight knot in Eq. (2.27). In particular, it follows that
$\varphi_{Q}(X)=\frac{1}{Q}\log\left(\frac{T(X^{Q})+(1-X^{2Q})\sqrt{\Delta(X^{Q})}}{2X^{4Q}}\right)=\frac{1}{Q}\log\left(\frac{\tilde{a}(X^{Q};1)+\sqrt{\tilde{a}(X^{Q};1)^{2}-4}}{2X^{2Q}}\right)\,,$ (4.152)
which gives us another way to numerically evaluate the integer constants $c_{s}$, $s\in{\mathbb{Z}}_{>0}$ for the geometric branch, following Eq. (3.86) for any choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime, in full agreement with our previous computations. In principle, one could use the formula for $\varphi_{Q}(X)$ with $P=Q=1$, that is,
@@
to derive the conjectural expression in Eq. (4.145), but we have not attempted to do so. The same formalism can be applied to the conjugate branch of the figure-eight knot, in which case the appropriate functions $s_{0,1}(x)$ are written in Eq. (4.138b) and the sign ambiguity in Eq. (3.4.1) is resolved by imposing the condition in Eq. (3.109) with $\alpha=\text{conj}$, which implies a choice of minus sign. Again, the corresponding computation of the integer constants $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, for the conjugate branch agrees with the results obtained above.
@@ 4.2 The three-twist knot
The state integral of the three-twist knot $\chi_{{\bf{5}}_{2}}(u;\hbar)$ can be expressed explicitly as bdp ; dimofte-levelk ; ggm2
$\chi_{{\bf{5}}_{2}}(u;\hbar)=$ ${\rm e}^{-\frac{\pi{\rm i}}{12}-\frac{5\pi{\rm i}}{3}c_{\mathsf{b}}^{2}}\Big{(}{\rm e}^{\pi{\rm i}(-2u^{2}-4c_{\mathsf{b}}u)}{\cal H}({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{4\pi{\mathsf{b}}u},{\rm e}^{4\pi{\mathsf{b}}u},;q){\cal H}({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{4\pi{\mathsf{b}}^{-1}u},{\rm e}^{4\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1})$ (4.154)
$+{\rm e}^{\pi{\rm i}(-2u^{2}+4c_{\mathsf{b}}u)}{\cal H}({\rm e}^{-2\pi{\mathsf{b}}u},{\rm e}^{-4\pi{\mathsf{b}}u},{\rm e}^{-4\pi{\mathsf{b}}u};q){\cal H}({\rm e}^{-2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-4\pi{\mathsf{b}}^{-1}u},{\rm e}^{-4\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1})$
$+{\cal H}({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{-2\pi{\mathsf{b}}u},1;q){\cal H}({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-2\pi{\mathsf{b}}^{-1}u},1;{\tilde{q}}^{-1})\Big{)}\,,$
where $c_{{\mathsf{b}}}={\rm i}({\mathsf{b}}+{\mathsf{b}}^{-1})/2$ and the $q$-special function ${\cal H}(x,y,z;q)$ is defined by121212${\cal H}(x,y;q)$ is closely related to the $q$-hypergeometric function.
${\cal H}(x,y,z;q)=(xq;q)_{\infty}(yq;q)_{\infty}\sum_{n=0}^{\infty}\frac{q^{n(n+1)}z^{n}}{(q;q)_{n}(xq;q)_{n}(yq;q)_{n}}\,,$ (4.155)
while $(xq;q)_{\infty}$ and $(xq;q)_{n}$ denote the quantum dilogarithm and the $q$-shifted factorial, respectively. See Appendix C for their definitions. As for the case of the figure-eight knot in Section 4.1, using Eq. (4.154), we can identify
$\phi^{(\rm{geom}),\rm{WKB}}(X;q)$ $=\log{\cal H}(X,X^{2},X^{2};q)\,,$ (4.156a)
$\phi^{(\rm{conj}),\rm{WKB}}(X;q)$ $=\log{\cal H}(X^{-1},X^{-2},X^{-2};q)\,,$ (4.156b)
$\phi^{(\rm{self}),\rm{WKB}}(X;q)$ $=\log{\cal H}(X,X^{-1},1;q)\,,$ (4.156c)
together with $C_{\rm{geom}}=C_{\rm{conj}}=C_{\rm{self}}={\rm e}^{\frac{3\pi{\rm i}}{4}}$ and
${{\rm i}s^{(\rm{geom})}_{0}(x)\over\hbar}+s^{(\rm{geom})}_{1}(x)+{{\rm i}s^{(\rm{geom})}_{0}(x_{D})\over\hbar_{D}}+s^{(\rm{geom})}_{1}(x_{D})$ $={x^{2}\over{\rm i}\hbar}+{2\pi x\over\hbar}+x+{5\pi^{2}{\rm i}\over 6\hbar}+{5{\rm i}\hbar\over 24}\,,$ (4.157a)
@@
Indeed, the three terms in the RHS of Eq. (4.154) capture the contributions to the exact partition function coming from the geometric, conjugate, and self-conjugate branches, respectively. These are the only three non-abelian branches of the $\bm{5}_{2}$-knot, as described in Example 2.2. Moreover, as before, we solve Eqs. (4.157a), (4.157b), and (4.157c) order by order in $x$ and $\hbar$ and obtain that
$s^{(\rm{geom})}_{0}(x)$ $=\frac{5\pi^{2}}{6}-\frac{x^{2}}{2}\,,\quad s^{(\rm{geom})}_{1}(x)=x\,,$ (4.158a)
$s^{(\rm{conj})}_{0}(x)$ $=\frac{5\pi^{2}}{6}-\frac{x^{2}}{2}\,,\quad s^{(\rm{conj})}_{1}(x)=-x\,,$ (4.158b)
$s^{(\rm{self})}_{0}(x)$ $=\frac{5\pi^{2}}{6}\,,\quad s^{(\rm{self})}_{1}(x)=0\,.$ (4.158c)
@@
In this way, we have fully determined the perturbative wave functions $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ of the three-twist knot in the resummed form of Eq. (3.39). As before, we verify successfully that they are annihilated by the $\hat{A}$-operator in Eq. (2.37), as expected.
@@
Let us now test the integrality conjecture in Eq. (3.44) for all three non-abelian branches. Expanding the RHS of Eqs. (4.156a), (4.156b), and (4.156c) gives
$\phi^{(\rm{geom}),\rm{WKB}}(X;q)$ $=\sum_{k\geq 1}\frac{q^{k}(X^{k}+X^{2k})}{k(q^{k}-1)}+\log\left(\sum_{n\geq 0}\frac{q^{n(n+1)}X^{2n}}{\prod_{i=1}^{n}(1-q^{i})(1-q^{i}X)(1-q^{i}X^{2})}\right)\,,$ (4.159a)
$\phi^{(\rm{conj}),\rm{WKB}}(X;q)$ $=\sum_{k\geq 1}\frac{q^{k}(X^{-k}+X^{-2k})}{k(q^{k}-1)}+\log\left(\sum_{n\geq 0}\frac{q^{n(n+1)}X^{n}}{\prod_{i=1}^{n}(1-q^{i})(X-q^{i})(X^{2}-q^{i})}\right)\,,$ (4.159b)
$\phi^{(\rm{self}),\rm{WKB}}(X;q)$ $=\sum_{k\geq 1}\frac{q^{k}(X^{k}+X^{-k})}{k(q^{k}-1)}+\log\left(\sum_{n\geq 0}\frac{q^{n(n+1)}X^{n}}{\prod_{i=1}^{n}(1-q^{i})(1-q^{i}X)(X-q^{i})}\right)\,.$ (4.159c)
@@
Both terms on the RHS of Eq. (4.159a) for the geometric branch contribute with positive powers of $X$. For the conjugate branch, we note that the negative powers of $X$ arise entirely from the first term on the RHS of Eq. (4.159b), while a factor of $q^{\sigma_{{\rm conj},n}}$ with
@@
$\sigma_{{\rm conj},n}=\left\lfloor\frac{n^{2}}{4}\right\rfloor\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (4.160)
@@
appears in the denominator of the power series in $X$ obtained by expanding the second term. Let us stress that this negative power of $q$ arises from the factors [^13] Note that there are twice as many contributing factors of $q^{-1}$ here compared to the conjugate branch of the figure-eight knot, in agreement with the observed formulae for the functions $\sigma_{{\rm conj},n}$ in Eqs. (4.140), (4.160). $\prod_{i=1}^{n}(X-q^{i})^{-1}(X^{2}-q^{i})^{-1}$ in the summand inside the logarithm, and it is absent from both the $X^{-1}$-series for the conjugate branch and the $X$-series for the geometric branch. Finally, in the case of the self-conjugate branch, the first term on the RHS of Eq. (4.159c) supplies both positive and negative powers of $X$, while the second term only adds to the $X$-series after expansion. Here, a more complicated factor of $q^{\sigma_{{\rm self},n}}$ with
@@
$\sigma_{{\rm self},n}=\left\lfloor\frac{1}{24}\left(P\left(\left\lfloor\frac{2(n+2)}{3}\right\rfloor\right)^{2}-1\right)\right\rfloor\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (4.161)
@@
where $P(n)$ denotes the $n$-th prime number, occurs in the denominator of the power series in $X$ after summing up the contribution from the expansion of the second term on the RHS of Eq. (4.159c). This originates from the factors $\prod_{i=1}^{n}(X-q^{i})^{-1}$ in the summand inside the logarithm.
@@
As we have done for the $\bm{4}_{1}$-knot, after extracting the coefficients $a^{(\alpha)}_{\pm n}(q)$, $n\in{\mathbb{Z}}_{>0}$, from the series for $\phi^{(\alpha),\rm{WKB}}(X;q)$, we obtain the functions $D_{\pm s}^{(\alpha)}(q)$, $s\in{\mathbb{Z}}_{>0}$, by applying the formulae in Eqs. (3.46), (3.49). Again, these are non-trivial polynomials with integer coefficients. In particular, the first few of them are
$D^{({\rm geom})}_{1}(q)$ $=q\,,\quad D^{({\rm geom})}_{2}(q)=-q(-1+q)\,,\quad D^{({\rm geom})}_{3}(q)=D^{({\rm geom})}_{4}(q)=-q^{3}\,,$ (4.162)
$D^{({\rm geom})}_{5}(q)$ $=q^{4}(-1+q^{2})\,,\quad D^{({\rm geom})}_{6}(q)=q^{4}(-1+2q^{2}+q^{4})\,,$
$D^{({\rm geom})}_{7}(q)$ $=q^{5}(-1+q(1+q)^{2})\,,$
@@
$D^{({\rm conj})}_{1}(q)$ $=D^{({\rm conj})}_{2}(q)=-1\,,\quad D^{({\rm conj})}_{3}(q)=-2-q\,,\quad D^{({\rm conj})}_{4}(q)=-(1+q)(1+2q)\,,$ (4.163)
@@
$D^{({\rm conj})}_{6}(q)$ $=-(1+q)(1+2q+3q^{2}+6q^{3}+7q^{4}+2q^{5})\,$
$D^{({\rm conj})}_{7}(q)$ $=-2-2q-8q^{2}-11q^{3}-20q^{4}-23q^{5}-32q^{6}-19q^{7}-6q^{8}-q^{9}\,,$
@@
$D^{({\rm self})}_{1}(q)=$ $2q\,,\quad D^{({\rm self})}_{2}(q)=1+q^{2}\,,\quad D^{({\rm self})}_{3}(q)=1+q^{2}+2q^{4}\,,$ (4.164)
@@
$D^{({\rm conj})}_{-1}(q)=D^{({\rm conj})}_{-2}(q)$ $=q\,,\quad D^{({\rm conj})}_{-s}(q)=0\quad s\geq 3\,,$ (4.165)
$D^{({\rm self})}_{-1}(q)$ $=q\,,\quad D^{({\rm self})}_{-s}(q)=0\quad s\geq 2\,.$

### v2

@@ 4.1 The figure-eight knot
$D^{({\rm conj})}_{-1}(q)=q,\qquad D^{({\rm conj})}_{-s}(q)=0\qquad s\geq 2.$ (4.10)
We have verified the integrality structure numerically up to $s=20$. It is then straightforward to compute the integers $c^{(\alpha)}_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, using equations (3.27) and (3.28), which give
$\bigl\{c^{({\rm geom})}_{s}\bigr\}_{s\in{\mathbb{Z}}_{>0}}=\{1,5,10,21,51,122,295,725,1792,4455,11133,27930,\dots\},$ (4.11a)
$\bigl\{c^{({\rm conj})}_{s}\bigr\}_{s\in{\mathbb{Z}}_{>0}}=(-1)\{0,4,9,20,50,121,294,724,1791,4454,11132,27929,\dots\},$ (4.11b)
while $c^{({\rm conj})}_{-s}=1$ for all $s\in{\mathbb{Z}}_{>0}$. We point out that the integers in equation (4.11a) appear to match numerically the sequence [^13] This is the entry A132461 in the on-line encyclopedia of integer sequences.
$\sum_{k=0}^{\lfloor s/2\rfloor}\left({s-k\choose k}+{s-k-1\choose k-1}\right)^{2},\qquad s\in{\mathbb{Z}}_{>0}.$ (4.12)
@@
$c^{({\rm geom})}_{s}+c^{({\rm conj})}_{s}=c^{({\rm conj})}_{-s},\qquad s\in{\mathbb{Z}}_{>0}.$ (4.13)
@@
Let us now go back to the quantum $A$-polynomial for the $\bm{4}_{1}$-knot written in equation (2.32). The corresponding second-order $q$-difference equation is
$\hat{A}_{\bm{4}_{1}}\bigl(\hat{P},\hat{X},q\bigr)\chi^{(\alpha),{\rm WKB}}(u;\hbar)=0,$ (4.14)
where $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ is the resummed perturbative wave function in equation (3.49) associated with the branch $\alpha$. In the rational case of $q={\rm e}^{2\pi{\rm i}{P\over Q}}$ with $P,Q\in{\mathbb{Z}}_{>0}$ coprime, equation (4.14) can be solved in closed form by directly applying the results of Section 3.4.1. For simplicity, let us consider the classical solution $\alpha={\rm geom}$ and drop the label from our notation. After suitably taking into account the exponential prefactor containing the functions $s_{0,1}(x)$ in equation (4.5a), the $q$-difference equation above assumes the form in equation (3.52) with the identifications
$A(X)$ $=q^{-2}X^{-2}{{C_{1}(X,q)\over C_{2}(X,q)}}=-q^{-2}X^{-2}{X^{2}-1\over q^{-1}X^{2}-1}\tilde{a}(X;q),$ (4.15a)
$B(X)$ $=q^{-2}X^{-4}{{C_{0}(X,q)\over C_{2}(X,q)}}=q^{-2}X^{-4}{qX^{2}-1\over q^{-1}X^{2}-1},$ (4.15b)
where $C_{0}$, $C_{1}$, $C_{2}$ are the coefficient functions defined in equation (2.33) and we have introduced
@@
Substituting the expressions in equations (4.15a) and (4.15b) into equation (3.61), we compute the matrix ${\cal M}_{Q}(X)$ and obtain explicit formulae for $\operatorname{tr}{\cal M}_{Q}(X)$, $\det{\cal M}_{Q}(X)$, and $\Delta_{Q}(X)$. Specifically, we find that
$\operatorname{tr}{\cal M}_{Q}(X)=X^{-4Q}\bigl(1-X^{Q}-2X^{2Q}-X^{3Q}+X^{4Q}\bigr),$ (4.17a)
@@
$\Delta_{Q}(X)=X^{-8Q}\bigl(1-X^{2Q}\bigr)^{2}\bigl(1-2X^{Q}-X^{2Q}-2X^{3Q}+X^{4Q}\bigr),$ (4.17c)
and plugging these into equation (3.72) with a choice of plus sign to satisfy the requirement in equation (3.71) for $\alpha={\rm geom}$, we obtain the functions $\Upsilon(X)$ and $\varphi_{Q}(X)$. Observe that all functions in equation (4.17), and thus also $\varphi_{Q}(X)$, depend on $X$ through $X^{Q}$, as expected, and we can identify
$\operatorname{tr}{\cal M}_{Q}(X)=X^{-4Q}T\bigl(X^{Q}\bigr)=X^{-2Q}\tilde{a}\bigl(X^{Q};1\bigr),$ (4.18a)
$\Delta_{Q}(X)=X^{-8Q}\bigl(1-X^{2Q}\bigr)^{2}\Delta\bigl(X^{Q}\bigr)=X^{-4Q}\bigl(\tilde{a}\bigl(X^{Q};1\bigr)^{2}-4\bigr),$ (4.18b)
where $\tilde{a}(X;q)$ is defined in equation (4.16), while $T(X)$ and $\Delta(X)$ are the functions introduced in equations (2.29a) and (2.29b), respectively, which appear in the formula for the solutions to the classical $A$-polynomial of the figure-eight knot in equation (2.28). In particular, it follows that
$\varphi_{Q}(X)$ $={{1\over Q}}\log\left({{T\bigl(X^{Q}\bigr)+\bigl(1-X^{2Q}\bigr)\sqrt{\Delta\bigl(X^{Q}\bigr)}\over 2X^{4Q}}}\right)$
$={{1\over Q}}\log\left({{\tilde{a}\bigl(X^{Q};1\bigr)+\sqrt{\tilde{a}\bigl(X^{Q};1\bigr)^{2}-4}\over 2X^{2Q}}}\right),$ (4.19)
which gives us another way to numerically evaluate the integer constants $c_{s}$, $s\in{\mathbb{Z}}_{>0}$, for the geometric branch, following equation (3.48) for any choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime. This is in full agreement with our previous computations. In principle, one could use the formula for $\varphi_{Q}(X)$ with $P=Q=1$, that is,
@@
to derive the conjectural expression in equation (4.12), but we have not attempted to do so. The same formalism can be applied to the conjugate branch of the figure-eight knot, in which case the appropriate functions $s_{0,1}(x)$ are written in equation (4.5b) and the sign ambiguity in equation (3.72) is resolved by imposing the condition in equation (3.71) with $\alpha=\mathrm{conj}$, which implies a choice of minus sign. Again, the corresponding computation of the integer constants $c_{\pm s}$$s\in{\mathbb{Z}}_{>0}$, for the conjugate branch agrees with the results obtained above.
@@ 4.2 The three-twist knot
The state integral of the three-twist knot $\chi_{{\bf{5}}_{2}}(u;\hbar)$ can be expressed explicitly as [5, 8, 17]
$\chi_{{\bf{5}}_{2}}(u;\hbar)={}$ ${\rm e}^{-{{\pi{\rm i}\over 12}}-{{5\pi{\rm i}\over 3}}c_{{\mathsf{b}}}^{2}}\bigl({\rm e}^{\pi{\rm i}(-2u^{2}-4c_{{\mathsf{b}}}u)}{\cal H}\bigl({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{4\pi{\mathsf{b}}u},{\rm e}^{4\pi{\mathsf{b}}u},;q\bigr)$
$\times{\cal H}\bigl({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{4\pi{\mathsf{b}}^{-1}u},{\rm e}^{4\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1}\bigr)$
$\phantom{\times}{}+{\rm e}^{\pi{\rm i}(-2u^{2}+4c_{{\mathsf{b}}}u)}{\cal H}\bigl({\rm e}^{-2\pi{\mathsf{b}}u},{\rm e}^{-4\pi{\mathsf{b}}u},{\rm e}^{-4\pi{\mathsf{b}}u};q\bigr)$
$\times{\cal H}\bigl({\rm e}^{-2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-4\pi{\mathsf{b}}^{-1}u},{\rm e}^{-4\pi{\mathsf{b}}^{-1}u};{\tilde{q}}^{-1}\bigr)$
$\phantom{\times}{}+{\cal H}\bigl({\rm e}^{2\pi{\mathsf{b}}u},{\rm e}^{-2\pi{\mathsf{b}}u},1;q\bigr){\cal H}\bigl({\rm e}^{2\pi{\mathsf{b}}^{-1}u},{\rm e}^{-2\pi{\mathsf{b}}^{-1}u},1;{\tilde{q}}^{-1}\bigr)\bigr),$ (4.21)
[…]
${\cal H}(x,y,z;q)=(xq;q)_{\infty}(yq;q)_{\infty}\sum_{n=0}^{\infty}{{q^{n(n+1)}z^{n}\over(q;q)_{n}(xq;q)_{n}(yq;q)_{n}}},$ (4.22)
while $(xq;q)_{\infty}$ and $(xq;q)_{n}$ denote the quantum dilogarithm and the $q$-shifted factorial, respectively. See Appendix C for their definitions. As for the case of the figure-eight knot in Section 4.1, using equation (4.21), we can identify
$\phi^{({\rm geom}),{\rm WKB}}(X;q)=\log{\cal H}\bigl(X,X^{2},X^{2};q\bigr),$ (4.23a)
$\phi^{({\rm conj}),{\rm WKB}}(X;q)=\log{\cal H}\bigl(X^{-1},X^{-2},X^{-2};q\bigr),$ (4.23b)
$\phi^{({\rm self}),{\rm WKB}}(X;q)=\log{\cal H}\bigl(X,X^{-1},1;q\bigr),$ (4.23c)
together with $\mathsf{C}_{{\rm geom}}=\mathsf{C}_{{\rm conj}}=\mathsf{C}_{{\rm self}}={\rm e}^{{{3\pi{\rm i}\over 4}}}$ and
${{\rm i}s^{({\rm geom})}_{0}(x)\over\hbar}+s^{({\rm geom})}_{1}(x)+{{\rm i}s^{({\rm geom})}_{0}(x_{D})\over\hbar_{D}}+s^{({\rm geom})}_{1}(x_{D})$
$\qquad={x^{2}\over{\rm i}\hbar}+{2\pi x\over\hbar}+x+{5\pi^{2}{\rm i}\over 6\hbar}+{5{\rm i}\hbar\over 24},$ (4.24a)
@@
Indeed, the three terms in the right-hand side of equation (4.21) capture the contributions to the exact partition function coming from the geometric, conjugate, and self-conjugate branches, respectively. These are the only three non-abelian branches of the $\bm{5}_{2}$-knot, as described in Example 2.2. Moreover, as before, we solve equations (4.24a), (4.24b), and (4.24c) order by order in $x$ and $\hbar$ and obtain that
$s^{({\rm geom})}_{0}(x)={{5\pi^{2}\over 6}}-{{x^{2}\over 2}},\qquad s^{({\rm geom})}_{1}(x)=x,$ (4.25a)
$s^{({\rm conj})}_{0}(x)={{5\pi^{2}\over 6}}-{{x^{2}\over 2}},\qquad s^{({\rm conj})}_{1}(x)=-x,$ (4.25b)
$s^{({\rm self})}_{0}(x)={{5\pi^{2}\over 6}},\qquad s^{({\rm self})}_{1}(x)=0.$ (4.25c)
@@
In this way, we have fully determined the perturbative wave functions $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ of the three-twist knot in the resummed form of equation (3.1). As before, we verify successfully that they are annihilated by the $\hat{A}-$operator in equation (2.38), as expected.
@@
Let us now test the integrality conjecture in equation (3.6) for all three non-abelian branches. Expanding the right-hand side of equations (4.23a), (4.23b), and (4.23c) gives
$\phi^{({\rm geom}),{\rm WKB}}(X;q)=\sum_{k\geq 1}{{q^{k}\bigl(X^{k}+X^{2k}\bigr)\over k\bigl(q^{k}-1\bigr)}}$
[…]
$\phi^{({\rm conj}),{\rm WKB}}(X;q)=\sum_{k\geq 1}{{q^{k}\bigl(X^{-k}+X^{-2k}\bigr)\over k\bigl(q^{k}-1\bigr)}}$
[…]
$\phi^{({\rm self}),{\rm WKB}}(X;q)=\sum_{k\geq 1}{{q^{k}\bigl(X^{k}+X^{-k}\bigr)\over k\bigl(q^{k}-1\bigr)}}$
[…]
@@
Both terms on the right-hand side of equation (4.26a) for the geometric branch contribute with positive powers of $X$. For the conjugate branch, we note that the negative powers of $X$ arise entirely from the first term on the right-hand side of equation (4.26b), while a factor of $q^{\sigma_{{\rm conj},n}}$ with
@@
$\sigma_{{\rm conj},n}=\left\lfloor{{n^{2}\over 4}}\right\rfloor,\qquad n\in{\mathbb{Z}}_{>0},$ (4.27)
@@
appears in the denominator of the power series in $X$ obtained by expanding the second term. Let us stress that this negative power of $q$ arises from the factors [^15] Note that there are twice as many contributing factors of $q^{-1}$ here compared to the conjugate branch of the figure-eight knot, in agreement with the observed formulae for the functions $\sigma_{{\rm conj},n}$ in equations (4.7) and (4.27). $\prod_{i=1}^{n}\bigl(X-q^{i}\bigr)^{-1}\bigl(X^{2}-q^{i}\bigr)^{-1}$ in the summand inside the logarithm, and it is absent from both the $X^{-1}$-series for the conjugate branch and the $X$-series for the geometric branch. Finally, in the case of the self-conjugate branch, the first term on the right-hand side of equation (4.26c) supplies both positive and negative powers of $X$, while the second term only adds to the $X$-series after expansion. Hence, a more complicated factor of $q^{\sigma_{{\rm self},n}}$ with
@@
$\sigma_{{\rm self},n}=\left\lfloor{{1\over 24}}\left(P\left(\left\lfloor{{2(n+2)\over 3}}\right\rfloor\right)^{2}-1\right)\right\rfloor,\qquad n\in{\mathbb{Z}}_{>0},$ (4.28)
@@
where $P(n)$ denotes the $n$-th prime number, occurs in the denominator of the power series in $X$ after summing up the contribution from the expansion of the second term on the right-hand side of equation (4.26c). This originates from the factors $\prod_{i=1}^{n}\bigl(X-q^{i}\bigr)^{-1}$ in the summand inside the logarithm.
@@
As we have done for the $\bm{4}_{1}$-knot, after extracting the coefficients $a^{(\alpha)}_{\pm n}(q)$, $n\in{\mathbb{Z}}_{>0}$, from the series for $\phi^{(\alpha),{\rm WKB}}(X;q)$, we obtain the functions $D_{\pm s}^{(\alpha)}(q)$, $s\in{\mathbb{Z}}_{>0}$, by applying the formulae in equations (3.8) and (3.11). Again, these are non-trivial polynomials with integer coefficients. In particular, the first few of them are
$D^{({\rm geom})}_{1}(q)=q,\qquad D^{({\rm geom})}_{2}(q)=-q(-1+q),\qquad D^{({\rm geom})}_{3}(q)=D^{({\rm geom})}_{4}(q)=-q^{3},$
$D^{({\rm geom})}_{5}(q)=q^{4}\bigl(-1+q^{2}\bigr),\qquad D^{({\rm geom})}_{6}(q)=q^{4}\bigl(-1+2q^{2}+q^{4}\bigr),$
$D^{({\rm geom})}_{7}(q)=q^{5}\bigl(-1+q(1+q)^{2}\bigr),$ (4.29)
@@
$D^{({\rm conj})}_{1}(q)=D^{({\rm conj})}_{2}(q)=-1,\qquad D^{({\rm conj})}_{3}(q)=-2-q,\qquad D^{({\rm conj})}_{4}(q)=-(1+q)(1+2q),$
@@
$D^{({\rm conj})}_{6}(q)=-(1+q)\bigl(1+2q+3q^{2}+6q^{3}+7q^{4}+2q^{5}\bigr),$
$D^{({\rm conj})}_{7}(q)=-2-2q-8q^{2}-11q^{3}-20q^{4}-23q^{5}-32q^{6}-19q^{7}-6q^{8}-q^{9},$ (4.30)
@@
$D^{({\rm self})}_{1}(q)=2q,\qquad D^{({\rm self})}_{2}(q)=1+q^{2},\qquad D^{({\rm self})}_{3}(q)=1+q^{2}+2q^{4},$
@@
$D^{({\rm conj})}_{-1}(q)=D^{({\rm conj})}_{-2}(q)=q,\qquad D^{({\rm conj})}_{-s}(q)=0,\qquad s\geq 3,$
$D^{({\rm self})}_{-1}(q)=q,\qquad D^{({\rm self})}_{-s}(q)=0,\qquad s\geq 2.$ (4.32)